# Reasoning Efficiently Through Adaptive Chain-of-Thought Compression: A Self-Optimizing Framework

### v1

@@ 6.1. Research Questions
  • RQ2: What is the contribution of each component of SEER? We conduct ablation studies by removing or varying parts of SEER, such as the CoT length filter $\lambda_{c}$ and the BoN parameter $\mathcal{N}$, to examine their impact.
  • RQ3: How well does SEER generalize across domains? We examine whether SEER remains effective when the fine-tuned data and test data come from different domains.
  • RQ4: Can SEER mitigate reasoning loops? We test whether SEER helps LLMs generate valid answers without falling into infinite reasoning loops.
@@ 6.2. RQ1: The Performance of SEER
In this subsection, we assess the performance of SEER in comparison with baseline methods. For the TokenSkip experiment, we set $\gamma=0.5$, meaning that we append the sequence “[EOS]0.5[EOS][EOS]0.5[EOS]” to the end of the input prompt. Similarly, for Naive BoN, we set $N=3$, which is consistent with the parameters used in evaluating our method. Table 2 reports task pass@1 accuracy ($\mathcal{A}$), average CoT length ($\lambda_{\text{avg}}$), and compression rate ($\rho$) under the 8K token constraint.
Methods MathQA-Python Defect-Detection Code-Search GSM-8K $\rho_{avg}(\%)$
$\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$
Base 63.7 1,456 - 44.7 1,836 - 72.4 472 - 90.1 1,309 - -
Self-Trained 74.6 1,177 19.2 49.3 1,055 42.5 72.5 468 0.9 89.2 792 39.5 25.5
Naive BoN (N=3) 73.6 1,191 18.2 49.8 1,042 43.3 74.9 478 -1.3 89.8 845 35.5 23.9
TokenSkip ($\gamma$=0.5) 1.6 1,196 17.9 36.2 2,421 -31.9 66.5 465 1.5 37.9 1,238 5.4 -1.8
TokenSkip ($\gamma$=0.8) 0.9 1,109 23.8 47.2 1,004 45.3 76.4 341 27.8 87.3 1,197 8.6 26.4
SEER 74.9 877 39.8 50.5 785 57.2 77.3 362 23.3 90.5 677 48.3 42.1
@@
    $\mathcal{A}(\%)$: Pass@1; $\lambda_{avg}$: Average output length (tokens);
    $\rho(\%)$: Compression rate compared to the Base model; $\rho_{avg}(\%)$: Average compression rate across all datasets.
@@
Table 2. The CoT Compression Performance Over Different Datasets (8K Token Limit)
Across all three software engineering tasks, SEER shows strong performance in both accuracy and CoT compression. On MathQA-Python, which requires code generation, it achieves 74.9% pass@1 with a 39.8% reduction in reasoning length. On Defect-Detection, it obtains the best pass@1 (50.5%) and the highest compression rate (57.2%). On Code-Search, it maintains a high pass@1 of 77.3% while compressing CoTs by 22.9%. Beyond software engineering, SEER also performs well on GSM-8K, a math reasoning dataset, reaching 90.5% pass@1 with a 48.3% reduction in CoT length. The accuracy gains are largely due to the mitigation of truncation and excessive reasoning, which often undermine model reliability. These results demonstrate that SEER consistently balances brevity and accuracy across diverse tasks.
[…]
Compared to TokenSkip, both Self-Trained and Naive BoN achieve more stable outcomes but still fall short of SEER. On Defect-Detection, their pass@1 remains below 50%, with compression rates under 30%. While they provide modest improvements over the base model, they cannot match SEER’s combination of high accuracy and substantial compression. This suggests that generic fine-tuning or simple BoN sampling is insufficient for the challenging trade-off between performance and efficiency.
[…]
Answer to RQ1: Compared with all baselines, SEER consistently achieves high accuracy and efficient CoT compression on both software engineering tasks and the math reasoning task. On average, SEER reduces CoT length by 42.1%, representing a 16.9% improvement over the best existing baseline, while preserving or even improving accuracy across diverse tasks.
6.3. RQ2: The Impact of Each Components
To investigate the contribution of each component, we conduct ablation studies on the MathQA-Python dataset by varying the BoN sampling size $\mathcal{N}$ and the CoT length filter threshold $\lambda_{c}$. The results are summarized in Table 3.
@@ 6.3. RQ2: The Impact of Each Components
Methods Settings Pass@1 $\mathcal{A}$ (%) Avg CoT Len $\lambda_{avg}$ (Tokens) Compression Rate $\rho$ (%)
@@
Without BoN $\lambda_{c}=2\times\lambda_{ada}$ 74.40 1,071 26.44
$\lambda_{c}=1.5\times\lambda_{ada}$ 74.19 918 36.95
$\lambda_{c}=\lambda_{ada}$ 73.50 878 39.70
$\lambda_{c}=0.5\times\lambda_{ada}$ 72.76 836 42.58
@@
When only BoN sampling is applied without the filter, increasing $\mathcal{N}$ from 1 to 3 improves accuracy from 73.49% to 74.45%, while the average CoT length decreases slightly from 1,191 to 1,177 tokens. However, further increasing $\mathcal{N}$ to 5 does not provide additional benefit and even slightly reduces accuracy (74.40%), indicating that the advantage of BoN sampling quickly saturates. In this setting, the compression rate remains relatively modest, ranging from 18.20% to 20.12%.
In contrast, applying only the adaptive filter leads to a substantial increase in compression. For instance, when $\lambda_{c}=2\times\lambda_{ada}$, the compression rate reaches 26.44%, and as the threshold becomes stricter ($\lambda_{c}=0.5\times\lambda_{ada}$), compression improves further to 42.58%. Nevertheless, overly aggressive filtering reduces accuracy, dropping from 74.40% to 72.76%. These results show that while the filter is more effective than BoN in reducing the redundancy in CoT, it risks discarding essential reasoning steps when applied too strongly.
To balance accuracy and compression rate, we set $\mathcal{N}=3$ and $\lambda_{c}=\lambda_{ada}$ in our method. When both components are combined in SEER with these settings, the method achieves the best trade-off, with the highest accuracy of 74.88% and a competitive compression rate of 39.77%. This demonstrates that the two modules are complementary: BoN sampling enhances robustness, while the adaptive filter enforces compactness, and together they enable SEER to deliver concise yet accurate reasoning.
@@
Answer to RQ2: The ablation results reveal that BoN sampling alone offers only limited gains, while the adaptive filter strongly enhances compression but may harm accuracy if too strict. Their combination achieves the best trade-off, yielding both higher accuracy and effective compression, highlighting the complementary roles of the two components.
6.4. RQ3: Generalizability of SEER
In this section, we evaluate the generalizability of SEER by fine-tuning it on one dataset and testing on the remaining three datasets. The results are presented in Table 4.
Table 4. The Generalizability Evaluation Result of SEER Across All Tasks
MathQA-Python Defect-Detection Code-Search GSM-8K
Models $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$
Base 63.7 1456 - 44.7 1,836 - 72.4 472 - 90.1 1,309 -
MathQA-Py Fine-tune 74.9 877 39.8 47.1 1131 38.4 77.3 395 16.3 87.2 748 42.9
Defect-Detection Fine-tune 64.2 1279 12.2 50.5 785 57.2 72.2 339 28.2 90.4 1128 13.8
Code-Search Fine-tune 64.1 1104 24.2 51.4 672 63.4 77.3 362 23.3 89.8 853 34.8
GSM-8K Fine-tune 62.5 1138 21.8 48.5 1483 19.2 79.6 435 7.8 90.5 677 48.3
Overall, the findings demonstrate that SEER exhibits strong cross-task generalization. Regardless of which dataset is used for fine-tuning, the resulting models generally achieve higher accuracy compared with the base model. Only a few cases show a slight decrease, such as fine-tuning on GSM-8K and then evaluating on MathQA-Python (62.5% vs. 63.7%). In all cases, the compression rate remains consistently above zero. For example, fine-tuning on MathQA-Python compresses average reasoning length on GSM-8K from 1,309 tokens to 748 tokens (42.9% compression), while still preserving high accuracy (87.2%). This indicates that SEER is able to reduce reasoning length while maintaining competitive accuracy even on out-of-domain tasks.
Interestingly, in some settings, the fine-tuned models even surpass those trained directly on their target dataset. For instance, the model fine-tuned on Code-Search achieves 51.4% accuracy on Defect-Detection, which is higher than the model fine-tuned on Defect-Detection (50.5%). These results suggest that SEER captures reusable reasoning patterns that extend beyond the dataset it was trained on, rather than overfitting to a single task.
Figure 4. Performance of SEER fine-tuned on different datasets when evaluated on HumanEval. Fine-tuning on coding-related datasets such as MathQA-Python and Code-Search yields the best generalization with shorter CoT lengths, while non-coding datasets like GSM-8K still provide moderate improvements over the base model.
To further assess its generalizability, we evaluate the four fine-tuned models on the HumanEval dataset, as shown in Figure 4. The results reveal two key insights. First, all fine-tuned models achieve noticeable improvements in both accuracy and CoT length over the base model, confirming that the knowledge learned from different tasks generalizes to program synthesis. Second, the reduction in reasoning length remains consistent, suggesting that SEER’s compression mechanism is stable even when applied to tasks outside the training domain. This demonstrates that the framework does not merely adapt to a specific dataset but instead learns reasoning strategies that generalize across diverse software engineering and mathematical tasks.
[…]
Answer to RQ3: SEER demonstrates strong cross-task generalization. Even when evaluated on out-of-domain datasets, it consistently improves accuracy while reducing CoT length. Fine-tuning on coding datasets further generalizes well to HumanEval, confirming that SEER learns reusable reasoning strategies beyond its training domain.
6.5. RQ4: The Ability of Reducing Reasoning Loops
Infinite reasoning loops occur when a model repeatedly generates identical reasoning segments until the output reaches the maximum context length and is truncated. This phenomenon wastes computational resources, inflates latency, and often produces incomplete or invalid final answers. Because such loops can arise unpredictably across models of different sizes, they represent a critical robustness concern. In this section, we investigate whether our approach can mitigate this problem.
For each dataset, we adapt our method to fine-tune a model specific to that dataset and evaluate it against the base model. The results in Table 5 show that SEER substantially reduces reasoning loops in all cases. On MathQA-Python, loop occurrences drop from 85 to 23 (72.9% reduction). On Code-Search, the count falls from 15 to 1 (93.3% reduction), effectively eliminating loops. The largest improvement appears in Defect-Detection, where loops decrease from 222 to just 7 (96.8% reduction). Even on GSM8K, which generally contains shorter reasoning chains, loops drop from 54 to 6 (88.9% reduction).
Table 5. The Performance of SEER in Reducing Truncation and Reasoning Loops
MathQA-Python Code-Search Defect-Detection GSM8K
Truncation Loops Truncation Loops Truncation Loops Truncation Loops
Base 110 85 18 15 235 222 59 54
SEER 30 23 1 1 7 7 8 6
@@ 6.5. RQ4: The Ability of Reducing Reasoning Loops
During our empirical analysis in Section 3, we find that most truncations in the base model are directly caused by such loops. This behavior wastes much of the available context budget and prevents the model from completing its reasoning. Our adaptive filtering strategy compresses CoT length, which mitigates loop formation and reduces truncation, as shorter reasoning provides fewer chances for repetitive output. This dual effect, loop suppression and truncation reduction, further enhances the robustness and practicality of our method in real-world deployments.
@@
[…]
Answer to RQ4: SEER effectively suppresses infinite reasoning loops across all datasets, with reductions ranging from 72.9% on MathQA-Python to 96.8% on Defect-Detection. Because most truncations in the base model stem from such loops, the adaptive filtering strategy not only prevents pathological repetition but also lowers truncation frequency.
@@ 7.1. The Impact of Training Approaches
An important aspect of our framework is its adaptability to different fine-tuning paradigms. Beyond full-parameter supervised fine-tuning (SFT), we also explored parameter-efficient alternatives such as LoRA. The comparison reveals two insights. First, while SFT unsurprisingly achieves the strongest overall performance, LoRA retains most of the benefits of our method, delivering substantial accuracy gains and compression improvements over the base model. This indicates that the effectiveness of adaptive CoT filtering does not depend on updating the entire parameter set.
@@
Pass@1 $\mathcal{A}$ (%) Avg CoT Len $\lambda_{avg}$ (Tokens) Compression Rate $\rho$ (%)
@@ 7.2.1. Internal Validity
Internal validity concerns whether the observed improvements of SEER truly arise from our proposed framework rather than confounding factors. One potential threat is the reliance on a single model family, namely the DeepSeek-R1-Distill-Qwen-7B. Although we carefully fine-tuned and evaluated under consistent settings to reduce bias, the conclusions may partially reflect properties of this architecture rather than general effectiveness. To mitigate this risk, we performed ablation studies (varying BoN sampling and adaptive filtering thresholds) to confirm that the observed gains stem from the specific design choices of SEER rather than incidental artifacts. Future work will extend the evaluation to other model families, including implicit CoT models that do not explicitly output reasoning traces. Moreover, testing on closed-source models with advanced reasoning capabilities (e.g., OpenAI O1, Gemini, Claude’s Thinking series) will help validate whether the improvements are architecture-independent.

### v2

@@ 6.1. Research Questions
  • RQ2: How well does SEER generalize across domains? We examine whether SEER remains effective when the fine-tuned data and test data come from different domains.
  • RQ3: What is the contribution of each component of SEER? We conduct ablation studies by removing or varying parts of SEER, such as the length-filter strictness $\alpha$ (which determines $\lambda_{c}$) and the BoN parameter $\mathcal{N}$, to examine their impact.
  • RQ4: How can SEER mitigate reasoning loops? We test whether SEER helps LLMs generate valid answers without encounter looping problem.
@@ 6.2. RQ1: The Performance of SEER
In this subsection, we evaluate the effectiveness of SEER under the 16K token constraint by comparing it with baselines. For baselines with tunable parameters, we adopt commonly used and representative settings. Specifically, TokenSkip is evaluated with $\gamma=0.5$ (i.e., compression ratio) to examine its maximum compression behavior, and with $\gamma=0.8$ to reflect a more conservative and widely used configuration. For Naive BoN, we sample $N=3$ candidates, which matches the sampling budget used in SEER. For BoN section in SEER, we set the same parameter, i.e., $N=3$. The results are summarized in Table 4.
Methods MathQA-Python Defect-Detection Code-Search $\rho_{avg}(\%)$
$pass@1(\%)$ $\lambda_{avg}$ $\rho(\%)$ $pass@1(\%)$ $\lambda_{avg}$ $\rho(\%)$ $pass@1(\%)$ $\lambda_{avg}$ $\rho(\%)$
Base 63.7 1,456 - 44.7 1,836 - 72.4 472 - -
TokenSkip ($\gamma$=0.5) 1.6 1,196 17.9 36.2 2,421 -31.9 66.5 465 1.5 -4.2
TokenSkip ($\gamma$=0.8) 0.9 1,109 23.8 47.2 1,004 45.3 76.4 362 23.3 30.8
Self-Trained 74.6 1,177 19.2 49.3 1,055 42.5 72.5 468 0.9 20.9
Naive BoN (N=3) 73.6 1,191 18.2 49.8 1,042 43.3 74.9 478 -1.3 20.1
Short CoT 55.6 1,236 15.1 46.6 1,198 34.8 69.7 398 15.7 21.9
SEER 74.9 877 39.8 50.5 785 57.2 77.3 341 27.8 41.6
@@
    $\lambda_{avg}$: Average output length (tokens);
    $\rho(\%)$: Compression rate compared to Base; $\rho_{avg}(\%)$: Average compression rate across all datasets;
    $\gamma$ in TokenSkip: compression ratio, lower $\gamma$ means higher compression rate.
@@
Table 4. The CoT Compression Performance Over Different Datasets (16K Token Limit)
Overall, SEER achieves the best performance in both reasoning efficiency and task accuracy across all three software engineering benchmarks. On MathQA-Python, SEER attains a pass@1 of 74.9% while reducing the average CoT length by 39.8%, indicating that substantial compression can be achieved without sacrificing correctness in code generation tasks. On Defect-Detection, SEER delivers the best overall performance, achieving the highest pass@1 (50.5%) together with the most aggressive compression (57.2%). Similarly, on Code-Search, SEER preserves the highest accuracy (77.3%) while shortening reasoning traces by 27.8%. These results demonstrate that SEER can effectively control CoT verbosity while maintaining, or even improving, task performance across diverse settings.
In contrast, TokenSkip exhibits unstable behavior across compression settings. Under a low $\gamma$ setting (i.e., high compression rate), it performs particularly poorly: on Defect-Detection, the average CoT length even increases compared to the base model, leading to more truncation and lower accuracy. Although a higher $\gamma$ improves stability, TokenSkip still underperforms SEER. Moreover, on MathQA-Python, its pass@1 drops to around 1% regardless of the compression setting, as its token-level compression ignores code structure and produces unexecutable code.
Other baselines that use fine-tuning methods without explicit compression operations, including Self-Trained and Naive BoN, show more stable behavior but offer limited gains. While these approaches achieve moderate improvements over the base model, their compression rates remain relatively low, and their accuracy consistently lags behind SEER. This suggests that generic fine-tuning or simple sampling strategies are insufficient to address the trade-off between reasoning length and performance.
Prompting-based methods also achieve a certain degree of CoT length reduction, with an average compression rate of 21.9% across the three datasets. However, this approach comes at a noticeable cost to accuracy, as concise prompt may suppress necessary intermediate reasoning. On the coding task (i.e., MathQA-Python), pass@1 drops substantially by 8.1%. These results further reinforce our Empirical Finding 3.2.
    Answer to RQ1: SEER consistently outperforms all baseline methods by achieving both high accuracy and substantial CoT compression across software engineering tasks. On average, SEER reduces CoT length by 41.6%, exceeding the strongest baseline by 10.8%, while improving pass@1 accuracy under strict context constraints.
6.3. RQ2: Generalization Performance of SEER
In this section, we evaluate the generalization capability of SEER across heterogeneous training domains. Specifically, we fine-tune the model on three different datasets and evaluate its performance on two unseen benchmarks, HumanEval and MBPP, as illustrated in Figure 4.
Figure 4. Performance of SEER fine-tuned on different software engineering datasets when evaluated on HumanEval and MBPP. Across all software engineering fine-tuning settings, SEER consistently improves model capability by achieving higher accuracy while reducing the average length of CoT reasoning, compared to the base model.
Figure 4 shows that fine-tuning SEER on different software engineering tasks consistently improves both task accuracy and reasoning efficiency when evaluated on unseen benchmarks. On HumanEval, the base model achieves about 78.0% accuracy with an average CoT length of 4,100 tokens; the fine-tuned models improve accuracy by 4.9 to 9.8 points (e.g., from 78.0% to 87.8%) while reducing the average reasoning length to around 2,400 tokens ( $\sim$40% reduction). On MBPP, the fine-tuned models improve accuracy by about 2.2 points (e.g., from 73.5% to 75.7%) and reduce the reasoning length to 1,680 to 2,100 tokens (up to 32% reduction). Overall, SEER achieves a stable CoT-length reduction on both benchmarks, indicating that it learns transferable compression behaviors.
    Answer to RQ2: On unseen benchmarks, SEER consistently improves accuracy while reducing reasoning length. Accuracy gains reach up to 9.8% on HumanEval and about 2.2% on MBPP, with a reduction of roughly 30% to 40% in CoT length.
6.4. RQ3: The Impact of Each Components
In this section, we conduct ablation studies to isolate the contribution of each module in SEER. We use MathQA-Python for ablation because it represents one of the most classic and widely studied tasks in software engineering (i.e., code generation). We vary (1) the BoN sampling size $\mathcal{N}$ while disabling the length filter, and (2) the strictness parameter $\alpha$ in the MAD-based cutoff (which determines $\lambda_{c}$) while removing BoN. We additionally report the full SEER configuration that combines both components. All other training and evaluation settings are kept identical. The results are summarized in Table 5.
@@ 6.4. RQ3: The Impact of Each Components
Methods Settings Pass@1 (%) Avg CoT Len $\lambda_{avg}$ (Tokens) Compression Rate $\rho$ (%)
@@
Without BoN $\alpha=2$ 74.40 1,071 26.44
[…]
$\alpha=1$ 73.50 878 39.70
[…]
@@
In the Without Filter setting (Table 5), we vary the BoN sampling size $\mathcal{N}$. Increasing $\mathcal{N}$ from 1 to 3 improves accuracy from 73.49% to 74.45%, while the average CoT length decreases slightly from 1,191 to 1,177 tokens. However, further increasing $\mathcal{N}$ to 5 does not provide additional benefit and slightly reduces accuracy (74.40%), indicating that the effect of BoN quickly saturates. In this setting, the compression rate remains modest (18.20% to 20.12%).
In the Without BoN setting (Table 5), we vary the strictness parameter $\alpha$ in our MAD-based cutoff $\lambda_{\mathrm{mad}}(\alpha)=\tilde{\lambda}+\alpha\cdot\mathrm{MAD}$. As the filter becomes stricter (from $\alpha=2$ to $\alpha=0.5$), the compression rate increases from 26.44% to 42.58%, but accuracy drops from 74.40% to 72.76%. This shows a clear trade-off: stronger filtering yields higher compression but can remove useful reasoning information.
We use $\mathcal{N}=3$ and $\alpha=1$ as the default setting in SEER. With both components enabled, SEER achieves the best balance, reaching the highest accuracy (74.88%) with a competitive compression rate (39.77%).
    Answer to RQ3: BoN alone provides only modest gains (e.g., accuracy increases from 73.49% to 74.45% when $\mathcal{N}$ increases from 1 to 3, with compression of 18.20% to 20.12%). Making the length filter stricter increases compression (from 26.44% to 42.58%) but reduces accuracy (from 74.40% to 72.76%). Using both components together achieves the best balance, reaching 74.88% accuracy with 39.77% compression.
6.5. RQ4: Effectiveness in Mitigating Reasoning Loops
Infinite reasoning loops occur when a model repeatedly generates identical reasoning segments until the output reaches the maximum context length and is truncated. This phenomenon wastes computational resources, inflates latency, and often produces incomplete or invalid final answers. Since such loops can arise unpredictably across models of different sizes, they represent a critical robustness concern. In this section, we investigate whether our approach can mitigate this problem.
We evaluate SEER on all three datasets, consistent with RQ1. For each dataset, we fine-tune a dataset-specific model using our method and compare it against the base model. Table 6 summarizes the results. SEER substantially reduces reasoning loops in all cases, e.g., from 85 to 23 on MathQA-Python (72.9%), from 15 to 1 on Code-Search (93.3%), and from 222 to 7 on Defect-Detection (96.8%).
Table 6. The Effectiveness of SEER in Mitigating Truncation and Reasoning Loops
MathQA-Python (NN=1883) Code-Search (NN=9604) Defect-Detection (NN=2732)
Truncation # (%) Loops # (%) Truncation # (%) Loops # (%) Truncation # (%) Loops # (%)
Base 110 (5.8%) 85 (4.5%) 18 (0.19%) 15 (0.16%) 235 (8.6%) 222 (8.1%)
@@ 6.5. RQ4: Effectiveness in Mitigating Reasoning Loops
In our Empirical Findings 3.2, we find that most truncations in the base model are directly caused by such loops. This behavior wastes much of the available context budget and prevents the model from completing its reasoning. Our adaptive filtering strategy reduces CoT length, mitigates loop formation and reduces truncation, as shorter reasoning provides fewer chances for repetitive output.
@@
    Answer to RQ4: SEER effectively suppresses infinite reasoning loops across all datasets, with reductions ranging from 72.9% on MathQA-Python to 96.8% on Defect-Detection. Because most truncations in the base model stem from such loops, the adaptive filtering strategy not only prevents looping problems but also lowers truncation frequency.
@@ 7.1. The Impact of Training Approaches
An important aspect of our framework is its adaptability to different fine-tuning paradigms. Beyond full-parameter supervised fine-tuning (SFT), we also explore parameter-efficient alternatives such as LoRA. The comparison reveals two insights. First, while SFT unsurprisingly achieves the strongest overall performance, LoRA retains most of the benefits of our method, delivering substantial accuracy gains and compression improvements over the base model. This indicates that the effectiveness of adaptive CoT filtering does not depend on updating the entire parameter set.
@@
$pass@1(\%)$ Avg CoT Len $\lambda_{avg}$ (Tokens) Compression Rate $\rho$ (%)
@@ 7.2.1. Internal Validity
Internal validity concerns whether the observed improvements of SEER truly arise from the proposed framework rather than confounding factors. One potential threat is the reliance on a single model family, namely DeepSeek-R1-Distill-Qwen-7B. Although all methods are fine-tuned and evaluated under consistent settings, some effects may reflect properties of this architecture. To mitigate this risk, we conduct ablation studies by varying BoN sampling and adaptive filtering thresholds, confirming that the gains are attributable to the core design of SEER rather than incidental artifacts.
Another potential concern is filtering bias. Our filter may exclude some difficult instances that naturally require longer CoTs. This is partially intended: SEER does not aim to create reasoning ability from scratch, but to encourage an already reasoning-capable model to produce concise, non-redundant reasoning. Since extremely long CoTs are often dominated by overthinking or looping, filtering these traces helps the model learn efficient reasoning patterns. To avoid removing problems entirely, filtering is applied to CoT traces rather than problem instances via BoN sampling. More importantly, our method does not incur a pass@1 loss after filtering, suggesting that this filtering strategy helps the model learn to generate reasoning that is both concise and effective.